OpenStudy (anonymous):

Todd has 4 different long-sleeved T-shirts and 2 baseball hats. If he selects one of each without looking, which ratio represents the probability that he'll be wearing his favorite T-shirt and favorite hat? one-eighth one-sixth two-sixthts two-eighths

OpenStudy (hoblos):

we have 4 different T-shirts then the probability that he'll be wearing his favorite T-shirt P(T) = 1/4 we have 2 different baseball hats then the probability that he'll be wearing his favorite baseball hats P(B) = 1/2 since the two events are independent the probability that he'll be wearing his favorite T-shirt and favorite hat is P(T intersection B) = (1/4)*(1/2) = 1/8

OpenStudy (anonymous):

I got 1/8, too